OpenStudy (anonymous):

Find the exact value of sin(165 degrees)

OpenStudy (anonymous):

sin(165) =sin(180-15) =sin(15) =sin(45-30) =sin(45)cos(30)-cos(45)sin(30) substitute for these standard angles and u'll get it.